Circular No. 6437

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996al IN NGC 7689 R. Evans and R. Benton, Coonabarabran, N.S.W.; and S. Beaman,

Pimpama, Queensland, report their visual discovery on July 22.71 UT with the Australian National University 1-m telescope at Siding Spring of a supernova (mag ∼ 14.0) located ∼ 30′′ north of the center of NGC 7689. The object does not appear on previous ESO survey films or on CCD survey images. J. Pogson and S. Ryan, Anglo-Australian Observatory, observing with the 3.9-m Anglo-Australian Telescope, measured the position of SN 1996al as α = 23^{h}33^{m}16^{s}.1, δ = −54^{o}05′02′′ (equinox 2000.0); a red spectrogram taken with an échelle spectrograph (UCLES) revealed Hα in emission, with a width coarsely estimated at some 20.0 nm, consistent with expansion velocities associated with a supernova.

KS 1731–260 E. H. Morgan and D. A. Smith, Massachusetts Institute of Technology,

report: “RXTE observations of KS 1731–260 performed on July 14.10–14.29 UT show persistent emission of 250 mCrab and a single type-I burst with a peak emission of 1.5 Crab at July 14.18. The total burst duration is 30 s, with a FWHM of 6 s in the band 2–60 keV. During the peak of the burst, a periodic signal at 524 Hz is detected. This signal is only seen for two seconds, starting 2 s after the rising edge of the burst. The pulse fraction is 12 ± 2 percent. This 524-Hz signal is similar to the 363-Hz signal found only during bursts from 4U 1728–34 (IAUC 6387). No periodic or quasiperiodic signals are detected in the persistent emission or in the tail of the burst. We speculate that we have detected the rotation frequency of the neutron star, in which case the spin period is 1.91 ms.”

SUPERNOVA 1996ak IN NGC 5021 M. Cavagna, Sormano, reports the following position for SN 1996ak ob-

tained with the 0.5-m reflector (+ CCD) on July 15.894 UT by V. Giuliani and F. Manca: α = 13^{h}12^{m}04^{s}.42, δ = +46^{o}11′38′′.4 (equinox 2000.0).

COMET C/1996 E1 (NEAT) Visual m_{1} estimates: June 15.19 UT, 14.4 (C. S. Morris, Lockwood

Valley, CA, 0.51-m reflector); 16.20, 14.4 (Morris); July 19.95, 10.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 21.97, 11.5 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector).

1996 July 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green